# Beyond MMD: Evaluating Graph Generative Models with Geometric Deep Learning

### v1

@@ II-C Graph Generative Models (GGMs)
The early development of GGMs leveraging deep learning techniques was spearheaded by GraphVAE [simonovsky2018graphvaegenerationsmallgraphs], which laid the groundwork for subsequent advancements in the field. GGMs leverage deep learning to capture complex relationships and dependencies between nodes and edges. This allows them to model intricate structural properties and, ideally, generate graphs that closely reflect the topology and statistical characteristics of the original dataset. As a result, GGMs are a powerful tool for graph generation tasks.
More formally, the objective in graph generation [guo2022systematicsurveydeepgenerative] is to learn a generative model capable of producing graphs that closely resemble a given set of input graphs $\{x_{i}\}$. This involves modeling the underlying data distribution and ensuring the ability to sample realistic graphs. Specifically:
  • $p_{\operatorname{data}}(x)$: represents the true data distribution from which the graphs are sampled, i.e., $x_{i}\sim p_{\operatorname{data}}(x)$, where $\sim$ means sampled or drawn.
  • $p_{\operatorname{model}}(x;\theta)$: represents the generative model, parameterized by $\theta$, which is designed to approximate $p_{\operatorname{data}}(x)$.
@@
    Ensure that $p_{\operatorname{model}}(x;\theta)$ closely approximates $p_{\operatorname{data}}(x)$.
@@
    Enable efficient sampling from $p_{\operatorname{model}}(x;\theta)$ to generate realistic graphs.
@@
$p_{\operatorname{model}}(x;\theta)=\prod^{n}_{t=1}p_{\operatorname{model}}(x_{t}\mid x_{1},\ldots,x_{t-1};\theta)$ (2)
In our case, $x_{t}$ will be the $t$-th action of adding a node or adding an edge. However, as the complexity of the graph increases, the evaluation of the generative process becomes more challenging and must consider various structural aspects and properties, such as degree distribution, clustering coefficient, community structure, motif prevalence, and connectivity patterns, which are essential for accurately reflecting the characteristics of real-world graphs. While some approaches rely on visual inspection of the generated graphs, we argue that this is not a reliable method for evaluating the performance of GGMs. Most recent approaches rely on Maximum Mean Discrepancy (MMD) as a metric to assess the distribution of graph properties in the generated ensemble.
@@
Graph Recurrent Attention Network (GRAN) [liao2020efficientgraphgenerationgraph] is an auto-regressive graph generative model that leverages the attention mechanism and Graph Neural Networks (GNNs) to generate graphs sequentially, simulating a growth process. The model generates blocks of nodes and their connections to the existing graph at each step.
Let $G=(V,E)$ be an undirected, unweighted graph with node set $V$ and edge set $E$. GRAN models the distribution over graphs via their adjacency matrices $A^{\pi}$, where $\pi$ denotes a particular node ordering. Unlike traditional autoregressive approaches that sequentially generate each edge or node, GRAN adopts a block-wise decoding strategy. The adjacency matrix $A^{\pi}$ is generated row-wise in blocks of size $B$. Without loss of generality, we can assume that the graph is undirected and that the lower triangle $L^{\pi}$ is generated and $A$ is symmetric. If that’s not the case, it can be easily achieved by generating the upper triangle in the same way. The model generates one block of $B$ rows of $L^{\pi}$ at a time, thus the number of steps to generate a graph is given by $T=\left\lceil N/B\right\rceil$. The joint probability of connections in the graph $G$ is factorized as:
$p(L^{\pi})=\prod_{t=1}^{T}p(L_{b_{t}}^{\pi}\mid L_{b_{1}}^{\pi},\ldots,L_{b_{t-1}}^{\pi})$ (3)
At each generation step $t$, GRAN constructs an augmented graph called $G_{t}$, which includes the already generated nodes $\{b_{1},\ldots,b_{t-1}\}$ and the current target block $b_{t}$, aiming at predicting the connectivity of $b_{t}$. Each node is initialized with a learned embedding via:
$h^{0}_{b_{i}}=WL_{b_{i}}^{\pi}+b,$ (4)
while the embeddings for new nodes are set to zero. In the architecture, a fixed number $R$ (rounds) of GNN message-passing layers are then applied. In each layer, messages are computed as:
@@
$a_{ij}^{r}=\sigma(g(\tilde{h}^{r}_{i}-\tilde{h}_{j}^{r})),\tilde{h}_{i}^{r}=[h_{i}^{r},x_{i}]$ (6)
where $x_{i}$ encodes node types (old vs new), $f$ and $g$ are learnable neural networks, and $\sigma$ is the sigmoid activation. Updated node representations are computed using a Gated Recurrent Unit (GRU) as in [li2017gatedgraphsequenceneural]. GRAN utilizes a mixture of Bernoulli distributions to model the probability of generating edge in the block $L_{b_{t}}^{\pi}$ by using the final node embeddings $h_{i}^{R}$ for each node $i$. Formally we have:
$p(L_{b_{t}}^{\pi}\mid L_{b_{1}}^{\pi},\ldots,L_{b_{t-1}}^{\pi})=\sum_{k=1}^{K}\alpha_{k}\prod_{i\in b_{t}}\prod_{1\leq j\leq i}\theta_{k,i,j}$ (7)
where, $K$ is the number of mixture components; $\alpha_{k}$ are the mixture weights that influence the edge generation and are computed using an MLP applied to node embedding differences. Similarly, the edge probability $\theta_{k,i,j}$ is obtained via another MLP followed by a sigmoid activation, also based on the difference between node embeddings.
@@
Efficient and Degree-Guided Graph Generative model (EDGE) [chen2023efficientdegreeguidedgraphgeneration] is a diffusion-based graph generative model. The main activity of EDGE is to invert the diffusion process on the edges: it starts with an empty graph (without edges) and a de-noising model adds edges progressively until a realistic graph is created.
Let $A^{0}$ be a graph from the data, EDGE defines a forward Markov chain such as:
$q(A^{1:T}\mid A^{0})=\prod^{T}_{t=1}q(A^{t}\mid A^{t-1}),$ (8)
where each transition independently removes edges with a probability $\beta_{t}$, in particular:
$\operatorname{\mathbf{Bern}}(A^{t};(1-\beta_{t})A^{t-1}+\beta_{t}p)$ (9)
At each step $t$, each existing edge is deleted, ensuring that the forward process converges to the all-zero adjacency matrix (empty graph). Finally, the posterior of the forward transition can be computed as:
$q(A^{t-1}\mid A^{t},A^{0})=\frac{q(A^{t}\mid A^{t-1})q(A^{t-1}\mid A^{0})}{q(A^{t}\mid A^{0})}$ (10)
Going on, rather than modeling all the potential edges, EDGE selects a small subset of nodes and only predicts edges among them. The nodes in this subset are the active nodes. So, $s^{t}:=\mathbbm{1}[d^{t-1}_{i}\neq d^{t}_{i}]$ with $d^{t}=\operatorname{deg}(A^{t})$, thus $s^{t}$ flags exactly those nodes whose degree changed (edges were removed). At each step, the model first predicts which nodes change degree ($s^{t}$), then adds edges only between the active nodes indicated by $s^{t}$.
To further match degree-guided generation, EDGE treats node degree $d^{0}$ as a random variable: it samples degree $d^{0}\sim p_{\theta}(d^{0})$, and it generates $A^{0}$ conditioning on $d^{0}$. Formally, we have the following decomposition, where we sample arbitrary node degree $d^{0}$ with $p_{\theta}(d^{0})$ for generating a graph with the degree constraint:
$p_{\theta}(A^{0},d^{0})=p_{\theta}(d^{0})p_{\theta}(A^{0}\mid d^{0})$ (11)
The denoising model will be:
$p_{\theta}(A^{0:T},s^{1:T},d^{0})=p_{\theta}(d^{0})p_{\theta}(A^{0:T},s^{1:T}\mid d^{0})$ (12)
The degree guidance is imposed by masking node selection so that nodes whose current degree $d_{i}^{t}$ has already reached the sampled target $d^{0}_{i}$ cannot become active.
By combining the edge-removal diffusion, active-node selection and degree conditioning, EDGE can scale discrete diffusion graph generation to thousands of nodes while matching key global graph statistics.
@@ III-A Evaluation Metric: Maximum Mean Discrepancy
Maximum Mean Discrepancy (MMD) [gretton2008kernelmethodtwosampleproblem] [JMLR:v13:gretton12a] is the most commonly used metric for evaluating the generative process of Graph Generative Models (GGMs) in the literature. It is a distance-measure between distributions $P(X)$ and $Q(Y)$ that is meant to capture the similarity between a specific property of the original and the generated graphs, $X$ and $Y$ respectively. Specifically, once a feature-extractor function $f$ is defined — mapping the graphs into a feature space — MMD computes the distance between the mean embeddings of the two distributions in a reproducing kernel Hilbert space (RKHS). The kernel function $k$ is mathematically defined as $k(x_{i},x_{j})=\langle f(x_{i}),f(x_{j})\rangle$, where $f(x_{i})\in\mathcal{H}$. For instance, $f$ may extract the degree distribution of the graphs, the clustering coefficient, or other graph-level features, which are then compared using a kernel function $k$ to measure the similarity between the feature representations of graphs in the feature space. Formally, given $n$ samples $X=\{x_{1},\ldots,x_{n}\}\subseteq\mathcal{X}$ and $m$ samples $Y=\{y_{1},\ldots,y_{m}\}\subseteq\mathcal{X}$, the biased empirical estimate of the MMD between $X$ and $Y$ can be defined as [DBLP:journals/corr/abs-2106-01098]:
@@
where $k$ is a positive definite kernel function, which is a measure of similarity between two graphs. That is, MMD accounts for the inner similarity between all the graph pairs in both the dataset and the generated graphs, and also for the cross-similarity between the two sets of graphs. The lower the MMD, the better.
This definition can quantify how closely the generated graphs resemble the original ones as a whole, according to some descriptor $f$, but has significant limitations that make it unsuitable for evaluating GGMs. First and foremost, the effectiveness of MMD heavily relies on the choice of the descriptor $f$ and kernel $k$ functions. In addition, MMD primarily compares the average value of the extracted property across the two sets, and may fail to capture differences in the internal distribution of that property within individual graphs. Secondly, as graphs are inherently complex structures characterized by intricate topological properties, optimizing one specific property may lead to the degradation of others.
This leads to a combination of MMD values that are hardly comparable and may not reflect the overall quality of the generated graphs. Additionally, MMD is unbounded and sensitive to outliers, which can significantly affect the results. Moreover, evaluating a single graph poses an additional challenge, as MMD is fundamentally designed to compare distributions rather than individual instances. We refer the reader to [DBLP:journals/corr/abs-2106-01098] for a more detailed discussion on the limitations of MMD and its application to graph generative models.
@@ III-B Neural Graph Similarity
The core idea is to use a Siamese Network [48e4505b7f124ee89368447f32b5cf0e], a neural architecture that learns task-specific similarity by comparing (the representation of) input pairs, to compute a similarity score between generated and original graphs in the embedding space. This allows us to assess how closely their learned representations align. In fact, the generation process should be able to produce graphs that resemble the original ones in some way, and a non-trivial similarity function can help quantify this resemblance that can hardly be captured by MMD or other metrics, as discussed above. The model could be trained, as common in deep learning, in an end-to-end fashion, using a set of labeled graphs, where the labels indicate whether the two graphs are similar or not.
Figure 1: Graph similarity using our proposed method, RGM. RGM leverages a Siamese architecture, where a Graph Neural Network (GNN) is trained to compute similarity between graphs. Input graphs are first pre-processed to extract informative node features (e.g., degree, $k$-core, local clustering coefficient, and some local statistics), then passed through the GNN to obtain embeddings. During training, the model receives triplets of graphs: an anchor, a positive (same class), and a negative (different class). Their embeddings (e.g., $\mathbf{h}_{a}$, $\mathbf{h}_{p}$, and $\mathbf{h}_{n}$, respectively in the figure) are optimized using a triplet loss function that minimizes the distance between anchor and positive graphs, while maximizing the distance to the negative one. This process encourages the model to learn a similarity function that maps graphs into an embedding space where structurally similar graphs are close together and dissimilar ones are farther apart. After training, the model can evaluate generated graphs (e.g., $g$ in the figure) by comparing them to anchor graphs from the training set. It computes similarity scores and assigns the generated graph to the class most represented among its closest anchors.

### v2

@@ II-C Graph Generative Models (GGMs)
The early development of GGMs leveraging deep learning techniques was spearheaded by GraphVAE [40], which laid the groundwork for subsequent advancements in the field. GGMs leverage deep learning to capture complex relationships and dependencies between nodes and edges. This allows them to model intricate structural properties and, ideally, generate graphs that closely reflect the topology and statistical characteristics of the original dataset. As a result, GGMs are a powerful tool for graph generation tasks.
More formally, the objective in graph generation [17] is to learn a generative model capable of producing graphs that closely resemble a given set of input graphs $\{G^{(i)}\}$. This involves modeling the underlying data distribution and ensuring the ability to sample realistic graphs. Specifically:
  • $p_{\text{data}}(G)$: represents the true data distribution from which the graphs are sampled, i.e., $G^{(i)}\sim p_{\text{data}}(G)$, where $\sim$ means sampled or drawn.
  • $p_{\theta}(G)$: represents the generative model, parameterized by $\theta$, which is designed to approximate $p_{\text{data}}(G)$.
@@
    Ensure that $p_{\theta}(G)$ closely approximates $p_{\text{data}}(G)$.
@@
    Enable efficient sampling from $p_{\theta}(G)$ to generate realistic graphs.
@@
$p_{\theta}(G^{(i)})=\prod_{t=1}^{T}p_{\theta}(A^{(i)}_{t}\mid A^{(i)}_{1},\ldots,A^{(i)}_{t-1})$ (2)
where $T$ is the total number of generation steps, and $A^{(i)}_{t}$ represents the adjacency matrix of $G^{(i)}$ after step $t$ of the generation process, which may involve adding a node, adding an edge, or more generally modifying existing connections. However, as the complexity of the graph increases, the evaluation of the generative process becomes more challenging and must consider various structural aspects and properties, such as degree distribution, clustering coefficient, community structure, motif prevalence, and connectivity patterns, which are essential for accurately reflecting the characteristics of real-world graphs. While some approaches rely on visual inspection of the generated graphs, we argue that this is not a reliable method for evaluating the performance of GGMs. Most recent approaches rely on Maximum Mean Discrepancy (MMD) as a metric to assess the distribution of graph properties in the generated ensemble.
@@
Graph Recurrent Attention Network (GRAN) [30] is an auto-regressive graph generative model that leverages the attention mechanism and Graph Neural Networks (GNNs) to generate graphs sequentially, simulating a growth process. The model generates blocks of nodes and their connections to the existing graph at each step.
Let $G=(V,E)$ be an undirected, unweighted graph with node set $V$ and edge set $E$. GRAN models the distribution over graphs via their adjacency matrices $A^{\pi}$, where $\pi$ denotes a particular node ordering. Unlike traditional autoregressive approaches that sequentially generate each edge or node, GRAN adopts a block-wise decoding strategy. The adjacency matrix $A^{\pi}$ is generated row-wise in blocks of size $B$. Without loss of generality, we can assume that the graph is undirected and that the lower triangle $L^{\pi}$ is generated and $A$ is symmetric. If that’s not the case, it can be easily achieved by generating the upper triangle in the same way. The model generates one block of $B$ rows of $L^{\pi}$ at a time, thus the number of steps to generate a graph is given by $T=\left\lceil N/B\right\rceil$, where $N=|V|$. The joint probability of connections in the graph $G$ is factorized as:
$p_{\theta}(L^{\pi})=\prod_{t=1}^{T}p_{\theta}(L_{b_{t}}^{\pi}\mid L_{b_{1}}^{\pi},\ldots,L_{b_{t-1}}^{\pi})$ (3)
where $L_{b_{t}}^{\pi}$ denotes the $t$-th block of $B$ rows in the lower triangular adjacency matrix $L^{\pi}$. At each generation step $t$, GRAN constructs an augmented graph called $G_{t}$, which includes the already generated nodes $\{b_{1},\ldots,b_{t-1}\}$ and the current target block $b_{t}$, aiming at predicting the connectivity of $b_{t}$. Each existing node $i$ is initialized with a learned embedding based on its corresponding row in the partial adjacency matrix:
$h^{0}_{i}=WL_{i}^{\pi}+b,$ (4)
where $L_{i}^{\pi}$ is the $i$-th row of the lower triangular matrix generated so far, while embeddings for newly added nodes in block $b_{t}$ are initialized to zero. In the architecture, a fixed number $R$ (rounds) of GNN message-passing layers are then applied. In each layer, messages are computed as:
@@
$a_{ij}^{r}=\sigma(g(\tilde{h}^{r}_{i}-\tilde{h}_{j}^{r})),\tilde{h}_{i}^{r}=[h_{i}^{r},c_{i}]$ (6)
where $c_{i}$ is an encoding of node types (existing vs newly added), $f$ and $g$ are learnable neural networks, and $\sigma$ is the sigmoid activation. Updated node representations are computed using a Gated Recurrent Unit (GRU) as in [29]. GRAN utilizes a mixture of Bernoulli distributions to model the probability of generating edges in the block $L_{b_{t}}^{\pi}$ by using the final node embeddings $h_{i}^{R}$ for each node $i$. Formally we have:
$p_{\theta}(L_{b_{t}}^{\pi}\mid L_{b_{1}}^{\pi},\ldots,L_{b_{t-1}}^{\pi})=\sum_{\ell=1}^{K}\alpha_{\ell}\prod_{i\in b_{t}}\prod_{1\leq j\leq i}\rho_{\ell,i,j}$ (7)
where $K$ is the number of mixture components; $\alpha_{\ell}$ are the mixture weights that influence the edge generation and are computed using an MLP applied to node embedding differences. The term $\rho_{\ell,i,j}\in[0,1]$ denotes the probability that an edge exists between nodes $i$ and $j$ under mixture component $\ell$, and is obtained via an MLP followed by a sigmoid activation applied to the difference between node embeddings $h_{i}^{R}$ and $h_{j}^{R}$.
@@
Efficient and Degree-Guided Graph Generative model (EDGE) [8] is a diffusion-based graph generative model. Like other diffusion models, EDGE is trained by learning to reverse a forward noising process. During training, a forward Markov chain progressively corrupts graphs by removing edges; during generation, a learned denoising model inverts this process, starting from an empty graph and adding edges progressively until a realistic graph is produced.
Following the standard diffusion model convention, let $\mathbf{A}_{0}$ denote the adjacency matrix of the target (clean) graph, and $\mathbf{A}_{t}$ the adjacency matrix at diffusion step $t$. EDGE defines a forward Markov chain that progressively corrupts the graph:
$q(\mathbf{A}_{1:T}\mid\mathbf{A}_{0})=\prod_{t=1}^{T}q(\mathbf{A}_{t}\mid\mathbf{A}_{t-1}),$ (8)
where each transition independently removes edges with probability $\beta_{t}$:
$(\mathbf{A}_{t})_{ij}\sim\operatorname{Bern}\bigl((1-\beta_{t})(\mathbf{A}_{t-1})_{ij}\bigr)$ (9)
At each step $t$, existing edges are stochastically deleted, ensuring that the forward process converges to the all-zero adjacency matrix (empty graph) as $t\to T$. The posterior of the forward transition can be computed as:
$q(\mathbf{A}_{t-1}\mid\mathbf{A}_{t},\mathbf{A}_{0})=\frac{q(\mathbf{A}_{t}\mid\mathbf{A}_{t-1})\,q(\mathbf{A}_{t-1}\mid\mathbf{A}_{0})}{q(\mathbf{A}_{t}\mid\mathbf{A}_{0})}$ (10)
Rather than modeling all potential edges, EDGE selects a small subset of active nodes and only predicts edges among them. The active node indicator is defined as $s_{i}^{t}:=\mathbbm{1}[(\mathbf{d}_{t-1})_{i}\neq(\mathbf{d}_{t})_{i}]$, where $\mathbf{d}_{t}=\operatorname{deg}(\mathbf{A}_{t})$ is the degree sequence at step $t$. Thus, $s_{i}^{t}$ flags nodes whose degree changed due to edge removal. At each step, the model first predicts which nodes change degree, then adds edges only between the active nodes.
To enable degree-guided generation, EDGE treats the target degree sequence $\mathbf{d}_{0}$ as a random variable: it first samples $\mathbf{d}_{0}\sim p_{\theta}(\mathbf{d}_{0})$, then generates $\mathbf{A}_{0}$ conditioned on $\mathbf{d}_{0}$. Formally:
$p_{\theta}(\mathbf{A}_{0},\mathbf{d}_{0})=p_{\theta}(\mathbf{d}_{0})\,p_{\theta}(\mathbf{A}_{0}\mid\mathbf{d}_{0})$ (11)
The full denoising (reverse) model is:
$p_{\theta}(\mathbf{A}_{0:T},s^{1:T},\mathbf{d}_{0})=p_{\theta}(\mathbf{d}_{0})\,p_{\theta}(\mathbf{A}_{0:T},s^{1:T}\mid\mathbf{d}_{0})$ (12)
Degree guidance is enforced by masking node selection: nodes whose current degree $(\mathbf{d}_{t})_{i}$ has already reached the target $(\mathbf{d}_{0})_{i}$ cannot become active.
By combining edge-removal diffusion, active-node selection, and degree conditioning, EDGE can scale discrete diffusion graph generation to thousands of nodes while matching key global graph statistics.
@@ III-A Evaluation Metric: Maximum Mean Discrepancy
Maximum Mean Discrepancy (MMD) [16] [15] is the most commonly used metric for evaluating the generative process of Graph Generative Models (GGMs) in the literature. It is a distance-measure between distributions $P(X)$ and $Q(Y)$ that is meant to capture the similarity between a specific property of the original and the generated graphs, $X$ and $Y$ respectively. Specifically, once a feature-extractor function $\phi$ is defined — mapping the graphs into a feature space — MMD computes the distance between the mean embeddings of the two distributions in a reproducing kernel Hilbert space (RKHS). The kernel function $k$ is mathematically defined as $k(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle$, where $\phi(x)\in\mathcal{H}$. For instance, $\phi$ may extract the degree distribution of the graphs, the clustering coefficient, or other graph-level features, which are then compared using a kernel function $k$ to measure the similarity between the feature representations of graphs in the feature space. Formally, given $n$ samples $X=\{x_{1},\ldots,x_{n}\}\subseteq\mathcal{X}$ and $m$ samples $Y=\{y_{1},\ldots,y_{m}\}\subseteq\mathcal{X}$, the biased empirical estimate of the MMD between $X$ and $Y$ can be defined as [35]:
@@
where $k$ is a positive definite kernel function, which is a measure of similarity between two samples. That is, MMD accounts for the inner similarity between all the graph pairs in both the dataset and the generated graphs, and also for the cross-similarity between the two sets of graphs. The lower the MMD, the better.
This definition can quantify how closely the generated graphs resemble the original ones as a whole, according to some descriptor $\phi$, but has significant limitations that make it unsuitable for evaluating GGMs. First and foremost, the effectiveness of MMD heavily relies on the choice of the descriptor $\phi$ and kernel $k$ functions. In addition, MMD primarily compares the average value of the extracted property across the two sets, and may fail to capture differences in the internal distribution of that property within individual graphs. Secondly, as graphs are inherently complex structures characterized by intricate topological properties, optimizing one specific property may lead to the degradation of others.
This leads to a combination of MMD values that are hardly comparable and may not reflect the overall quality of the generated graphs. Additionally, MMD is unbounded and sensitive to outliers, which can significantly affect the results. Moreover, evaluating a single graph poses an additional challenge, as MMD is fundamentally designed to compare distributions rather than individual instances. We refer the reader to [35, shirzad2022evaluating] for a more detailed discussion on the limitations of MMD and its application to graph generative models.
@@ III-B Neural Graph Similarity
The core idea is to use a Siamese Network [3], a neural architecture that learns task-specific similarity by comparing (the representation of) input pairs, to compute a similarity score between generated and original graphs in the embedding space. This allows us to assess how closely their learned representations align. In fact, the generation process should be able to produce graphs that resemble the original ones in some way, and a non-trivial similarity function can help quantify this resemblance that can hardly be captured by MMD or other metrics, as discussed above. The model could be trained, as common in deep learning, in an end-to-end fashion, using a set of labeled graphs, where the labels indicate whether the two graphs are similar or not.
Figure 1: Graph similarity using our proposed method, RGM. RGM leverages a Siamese architecture, where a Graph Neural Network (GNN) — specifically a Graph Attention Network (GAT) — is trained to compute similarity between graphs. Input graphs are first pre-processed to extract informative node features (e.g., degree, $k$-core, local clustering coefficient, and chi-squared statistics), then passed through the GNN to obtain embeddings. During training, the model receives triplets of graphs: an anchor, a positive (same class), and a negative (different class). Their embeddings (e.g., $\mathbf{h}_{a}$, $\mathbf{h}_{p}$, and $\mathbf{h}_{n}$, respectively in the figure) are optimized using a triplet loss function that minimizes the distance between anchor and positive graphs, while maximizing the distance to the negative one. This process encourages the model to learn a similarity function that maps graphs into an embedding space where structurally similar graphs are close together and dissimilar ones are farther apart. After training, the model can evaluate generated graphs (e.g., $g$ in the figure) by comparing them to anchor graphs from the training set. It computes similarity scores and assigns the generated graph to the class most represented among its closest anchors.